# Residual-based Chebyshev filtered subspace iteration for sparse Hermitian eigenvalue problems tolerant to inexact matrix-vector products with applications to Kohn-Sham density functional theory

### v4

@@ 2.1 Chebyshev filtered subspace iteration and convergence properties
ChFSI leverages the properties of Chebyshev polynomials to efficiently filter out the components of the unwanted eigenvectors (corresponding to the remaining mnm-n largest eigenvalues), thus enriching the trial subspace with the desired eigenvectors. To this end, we define the Chebyshev polynomial of degree $k$ as $T_{k}(x)$ and note that these polynomials exhibit the fastest growth in their magnitude when $\lvert x\rvert\,>1$ while remaining bounded between [1,1][-1,1] when $\lvert x\rvert\,\leq 1$. To exploit this fast growth property of $T_{k}(x)$, an affine transformation that maps the largest mnm-n eigenvalues to [1,1][-1,1] is defined, and consequently, the desired smallest $n$ eigenvalues get mapped to values lying in $(-\infty,-1]$. To this end, we define the center of the unwanted spectrum as $c=(\lambda_{n+1}+\lambda_{m})/2$ and the half-width of the unwanted spectrum as $e=(\lambda_{m}-\lambda_{n+1})/2$ and hence the required affine transformation can now be represented as $\mathcal{L}(x)=(x-c)/e$. Standard implementations of ChFSI also scale the Chebyshev polynomials to prevent overflow [saadChebyshevAccelerationTechniques1984, saadNumericalMethodsLarge2011, zhouParallelSelfconsistentfieldCalculations2006, zhouSelfconsistentfieldCalculationsUsing2006, zhouChebyshevfilteredSubspaceIteration2014, winkelmannChASE2019] and consequently the scaled and shifted Chebyshev polynomials are defined as $C_{k}(x)=T_{k}(\mathcal{L}(x))/T_{k}(\mathcal{L}(a_{low}))$ where $a_{low}\leq\lambda_{1}$. We note that the ChFSI procedure for solving the eigenproblem in eq. 1 for the case of sparse matrices [^2] For dense matrices the common practice is to employ Cholesky factorization of B to convert the generalized eigenvalue problem to the standard eigenvalue problem. is usually devised with the matrix $\textbf{H}=\textbf{B}^{-1}\textbf{A}$ that has the same eigenpairs as $\textbf{A}\textbf{u}_{j}=\lambda_{j}\textbf{B}\textbf{u}_{j}$. This procedure for computing the smallest eigenpairs of eq. 1 up to a specified tolerance $\tau$ on the eigenproblem residual norm, is summarized in algorithm 1. We note that in algorithm 1 we do not explicitly perform an orthogonalization step for the filtered subspace and instead opt to perform the Rayleigh-Ritz step for a non-orthogonal basis, which ensures that the resulting Ritz vectors are B-orthonormal. This choice was made for both ease of analysis and computational efficiency.
@@
Initial Guess: Let $\textbf{X}^{(0)}=\begin{bmatrix}\textbf{x}_{1}^{(0)}&\textbf{x}_{2}^{(0)}&\dots&\textbf{x}_{n}^{(0)}\end{bmatrix}$ be the initial guess of the eigenvectors ($\{\textbf{u}_{j}\}$).
while $r_{j}^{(i+1)}=\lVert\textbf{A}\textbf{x}^{(i+1)}_{j}-\epsilon_{j}^{(i+1)}\textbf{B}\textbf{x}^{(i+1)}_{j}\rVert\,\geq\tau$ do
  Chebyshev Filtered Subspace Construction: Construct $\textbf{Y}_{p}^{(i)}=C_{p}(\textbf{H})\textbf{X}^{(i)}$ using the recurrence:
where $\textbf{Y}_{0}^{(i)}=\textbf{X}^{(i)}$ and $\textbf{Y}_{1}^{(i)}=\frac{\sigma_{1}}{e}(\textbf{H}-c\textbf{I})\textbf{X}^{(i)}$ and $\sigma_{k+1}=1/\left(\sigma_{k}-\frac{2}{\sigma_{1}}\right)$ with $\sigma_{1}=e/(a_{low}-c)$. Here $\textbf{Y}_{k}^{(i)}=C_{k}(\textbf{H})\textbf{X}^{(i)}$ for $k=0,1,\dots,p$ as described in algorithm 2
  Rayleigh-Ritz step: Solve the smaller $n\times n$ dense generalized eigenvalue problem. Denoting the transpose conjugate of a matrix by we have ${\textbf{Y}_{p}^{(i)}}^{\dagger}\textbf{A}\textbf{Y}_{p}^{(i)}\textbf{E}={\textbf{Y}_{p}^{(i)}}^{\dagger}\textbf{B}\textbf{Y}_{p}^{(i)}\textbf{E}\boldsymbol{\Lambda}$, where E is the eigenvector matrix and $\boldsymbol{\Lambda}$ is the diagonal matrix with the eigenvalues $\left\{\epsilon_{j}^{(i+1)}\right\}_{j=1}^{n}$ as its entries. The Ritz vectors given by $\textbf{X}^{(i+1)}=\textbf{Y}_{p}^{(i)}\textbf{E}$ are B-orthonormal while the Ritz-Values are given by $\boldsymbol{\Lambda}^{(i+1)}=\boldsymbol{\Lambda}$.
@@ 2.1.1 Convergence analysis of ChFSI
To ensure completeness, we now provide a mathematical justification for the convergence of ChFSI. Note that $\lVert\cdot\rVert\,$ implies the 2-norm for vectors and vector-induced matrix 2-norm for matrices (spectral norm) throughout this work. In order to analyze convergence, we first define the maximum principal angle between two subspaces [chenSpectralMethodsData2021, zhuAnglesSubspacesTheir2013, watkinsMatrixEigenvalueProblem2007].
@@
Let $\mathcal{X}\subset\mathbb{C}^{m\times m}$ and $\mathcal{Y}\subset\mathbb{C}^{m\times m}$ be subspaces of dimension $n$. The maximum principal angle between the two subspaces denoted by $\angle(\mathcal{X},\mathcal{Y})$ is defined as
$\sin\angle(\mathcal{X},\mathcal{Y})=\lVert\textbf{X}_{\perp}^{\dagger}\textbf{Y}\rVert\,=\lVert\textbf{Y}_{\perp}^{\dagger}\textbf{X}\rVert\,$
where $\textbf{X}\in\mathbb{C}^{m\times n}$ and $\textbf{Y}\in\mathbb{C}^{m\times n}$ are matrices whose columns form an orthonormal basis for $\mathcal{X}$ and $\mathcal{Y}$ respectively. $\textbf{X}_{\perp}\in\mathbb{C}^{m\times(m-n)}$ and $\textbf{Y}_{\perp}\in\mathbb{C}^{m\times(m-n)}$ are matrices whose columns form an orthonormal basis spanning the orthogonal complements of $\mathcal{X}$ and $\mathcal{Y}$ respectively.
@@
Since the maximum principal angle given by definition 2.1 is invariant under unitary transformations, we have ${\angle(\mathcal{S}^{(i+1)},\mathcal{S})}={\angle(\hat{\textbf{U}}^{\dagger}\mathcal{S}^{(i+1)},\hat{\textbf{U}}^{\dagger}\mathcal{S})}$ and also ${\angle(\mathcal{S}^{(i)},\mathcal{S})}={\angle(\hat{\textbf{U}}^{\dagger}\mathcal{S}^{(i)},\hat{\textbf{U}}^{\dagger}\mathcal{S})}$. Since the space defined by $\hat{\textbf{U}}^{\dagger}\mathcal{S}$ is the same as $\mathcal{E}=\mathcal{R}(\textbf{E})$, we have, ${\angle(\mathcal{S}^{(i+1)},\mathcal{S})}={\angle(\hat{\textbf{U}}^{\dagger}\mathcal{S}^{(i+1)},\mathcal{E})}$ and ${\angle(\mathcal{S}^{(i)},\mathcal{S})}={\angle(\hat{\textbf{U}}^{\dagger}\mathcal{S}^{(i)},\mathcal{E})}$. Consider a partitioning of the matrix $\hat{\textbf{U}}^{\dagger}\hat{\textbf{X}}^{(i)}$ as $\hat{\textbf{U}}^{\dagger}\hat{\textbf{X}}^{(i)}=\Bigl[\begin{smallmatrix}\hat{\textbf{Z}}_{1}^{(i)}\\ \hat{\textbf{Z}}_{2}^{(i)}\end{smallmatrix}\Bigr]$ where $\hat{\textbf{Z}}_{1}^{(i)}=\hat{\textbf{U}}_{1}^{\dagger}\hat{\textbf{X}}^{(i)}$ and $\hat{\textbf{Z}}_{2}^{(i)}=\hat{\textbf{U}}_{2}^{\dagger}\hat{\textbf{X}}^{(i)}$. We note that the assumption $\mathcal{S}^{(i)}\cap\mathcal{S}_{\perp}=\left\{0\right\}$ ensures that $\hat{\textbf{Z}}_{1}^{(i)}$ is invertible, and consequently, we can write
[…]
Subsequently from lemma A.1 we have $\tan{\angle(\mathcal{S}^{(i)},\mathcal{S})}=\lVert\hat{\textbf{Z}}_{2}^{(i)}{{}\hat{\textbf{Z}}_{1}^{(i)}}^{-1}\rVert\,$. Further, partitioning the diagonal matrix of eigenvalues $\boldsymbol{\Lambda}$, we have $\boldsymbol{\Lambda}=\Bigl[\begin{smallmatrix}\boldsymbol{\Lambda}_{1}&0\\ 0&\boldsymbol{\Lambda}_{2}\end{smallmatrix}\Bigr]$ where $\boldsymbol{\Lambda}_{1}$ is the $n\times n$ diagonal matrix comprising of the wanted eigenvalues and $\boldsymbol{\Lambda}_{2}$ is the $(m-n)\times(m-n)$ diagonal matrix comprising of the unwanted eigenvalues. We can now write
[…]
From lemma A.1 we now have $\tan{\angle(\mathcal{S}^{(i+1)},\mathcal{S})}=\lVert C_{p}(\boldsymbol{\Lambda}_{2})\hat{\textbf{Z}}_{2}^{(i)}{{}\hat{\textbf{Z}}_{1}^{(i)}}^{-1}C_{p}(\boldsymbol{\Lambda}_{1})^{-1}\rVert\,$, allowing us to write
$\tan{\angle(\mathcal{S}^{(i+1)},\mathcal{S})}\leq\lVert C_{p}(\boldsymbol{\Lambda}_{2})\rVert\,\lVert\hat{\textbf{Z}}_{2}^{(i)}{{}\hat{\textbf{Z}}_{1}^{(i)}}^{-1}\rVert\,\lVert C_{p}(\boldsymbol{\Lambda}_{1})^{-1}\rVert\,=\left\lvert\frac{C_{p}(\lambda_{n+1})}{C_{p}(\lambda_{n})}\right\rvert\,\tan{\angle(\mathcal{S}^{(i)},\mathcal{S})}$
@@ 3 Approximate Chebyshev Filtered Subspace Construction
We note that the computationally dominant step in Chebyshev filtered subspace construction is the evaluation of the sparse-matrix multi-vector product $\textbf{H}\textbf{Y}_{k}^{(i)}$ in eq. 2. A straightforward way to accelerate this step is to use approximations in the computation of $\textbf{H}\textbf{Y}_{k}^{(i)}$, allowing for improved efficiency. This should, in principle, allow for the use of various efficient approximate matrix multiplication techniques, including but not limited to mixed-precision arithmetic. However, we note that this requires an understanding of the convergence properties of ChFSI when such approximations are employed. We now adapt theorem 2.2 for the case where approximations are used in computing $\textbf{H}\textbf{Y}_{k}^{(i)}$.

### v5

@@ 2.1 Chebyshev filtered subspace iteration and convergence properties
ChFSI leverages the properties of Chebyshev polynomials to efficiently filter out the components of the unwanted eigenvectors (corresponding to the remaining mnm-n largest eigenvalues), thus enriching the trial subspace with the desired eigenvectors. To this end, we define the Chebyshev polynomial of degree $k$ as $T_{k}(x)$ and note that these polynomials exhibit the fastest growth in their magnitude when $\lvert x\rvert\,>1$ while remaining bounded between [1,1][-1,1] when $\lvert x\rvert\,\leq 1$. To exploit this fast growth property of $T_{k}(x)$, an affine transformation that maps the largest mnm-n eigenvalues to [1,1][-1,1] is defined, and consequently, the desired smallest $n$ eigenvalues get mapped to values lying in $(-\infty,-1]$. To this end, we define the center of the unwanted spectrum as $c=(\lambda_{m}+\lambda_{n+1})/2$ and the half-width of the unwanted spectrum as $e=(\lambda_{m}-\lambda_{n+1})/2$ and hence the required affine transformation can now be represented as $\mathcal{L}(x)=(x-c)/e$. Standard implementations of ChFSI also scale the Chebyshev polynomials to prevent overflow [37, 38, 52, 53, 51, 47] and consequently the scaled and shifted Chebyshev polynomials are defined as $C_{k}(x)=T_{k}(\mathcal{L}(x))/T_{k}(\mathcal{L}(a_{low}))$ where $a_{low}\leq\lambda_{1}$. In practice, the exact eigenvalues are unknown, and the filter is parameterised by estimates: a lower bound $\lambda_{\min}\leq\lambda_{1}$ (taking the role of $a_{low}$), an upper bound $\lambda_{\max}\geq\lambda_{m}$ on the spectrum, and a threshold $\lambda_{T}$ separating the wanted from the unwanted eigenvalues (so that $\lambda_{T}\approx\lambda_{n+1}$). In terms of these estimates the affine-transformation parameters become $c=(\lambda_{\max}+\lambda_{T})/2$ and $e=(\lambda_{\max}-\lambda_{T})/2$. We note that the ChFSI procedure for solving the eigenproblem in eq. 1 for the case of sparse matrices [^2] For dense matrices the common practice is to employ Cholesky factorization of B to convert the generalized eigenvalue problem to the standard eigenvalue problem. is usually devised with the matrix $\textbf{H}=\textbf{B}^{-1}\textbf{A}$ that has the same eigenpairs as $\textbf{A}\textbf{u}_{j}=\lambda_{j}\textbf{B}\textbf{u}_{j}$. This procedure for computing the smallest eigenpairs of eq. 1 up to a specified tolerance $\tau$ on the eigenproblem residual norm, is summarized in algorithm 1. We note that in algorithm 1 we do not explicitly perform an orthogonalization step for the filtered subspace and instead opt to perform the Rayleigh-Ritz step for a non-orthogonal basis, which ensures that the resulting Ritz vectors are B-orthonormal. This choice was made for both ease of analysis and computational efficiency.
The key computational kernel of ChFSI is the Chebyshev polynomial filter, which constructs the filtered subspace $\textbf{Y}_{p}^{(i)}=C_{p}(\textbf{H})\textbf{X}^{(i)}$ via the three-term recurrence
where $\textbf{Y}_{0}^{(i)}=\textbf{X}^{(i)}$, $\textbf{Y}_{1}^{(i)}=\frac{\sigma_{1}}{e}(\textbf{H}-c\textbf{I})\textbf{X}^{(i)}$, and the scaling coefficients satisfy $\sigma_{k+1}=1/\left(\frac{2}{\sigma_{1}}-\sigma_{k}\right)$ with $\sigma_{1}=e/(a_{low}-c)$. Here $\textbf{Y}_{k}^{(i)}=C_{k}(\textbf{H})\textbf{X}^{(i)}$ for $k=0,1,\dots,p$. The filtering procedure is detailed in algorithm 2. Following the filtering step, a Rayleigh–Ritz projection is performed. Denoting the conjugate transpose by , one solves the reduced $n\times n$ dense generalized eigenvalue problem
${\textbf{Y}_{p}^{(i)}}^{\dagger}\textbf{A}\textbf{Y}_{p}^{(i)}\textbf{E}^{(i+1)}={\textbf{Y}_{p}^{(i)}}^{\dagger}\textbf{B}\textbf{Y}_{p}^{(i)}\textbf{E}^{(i+1)}\boldsymbol{\Lambda}^{(i+1)},$
where $\textbf{E}^{(i+1)}$ is the eigenvector matrix and $\boldsymbol{\Lambda}^{(i+1)}=\mathrm{diag}(\epsilon_{1}^{(i+1)},\dots,\epsilon_{n}^{(i+1)})$ contains the Ritz values. The updated Ritz vectors are $\textbf{X}^{(i+1)}=\textbf{Y}_{p}^{(i)}\textbf{E}^{(i+1)}$, which are B-orthonormal by construction. These two steps are repeated until the residual norm $r_{j}^{(i+1)}=\lVert\textbf{A}\textbf{x}_{j}^{(i+1)}-\epsilon_{j}^{(i+1)}\textbf{B}\textbf{x}_{j}^{(i+1)}\rVert\,$ falls below a desired tolerance $\tau$ for all $j=1,\dots,n$. The overall procedure is summarized in algorithm 1.
@@
INPUTS: Hermitian matrix $\textbf{A}\in\mathbb{C}^{m\times m}$, SPD matrix $\textbf{B}\in\mathbb{C}^{m\times m}$, number of wanted eigenpairs $n$, Chebyshev polynomial degree $p$, spectral bounds $\lambda_{\min},\lambda_{\max},\lambda_{T}$, residual tolerance $\tau$.
OUTPUT: Approximate eigenvectors $\textbf{X}^{(i+1)}\in\mathbb{C}^{m\times n}$ and eigenvalues $\boldsymbol{\Lambda}^{(i+1)}\in\mathbb{R}^{n\times n}$.
Step 0. Choose an initial guess $\textbf{X}^{(0)}=\begin{bmatrix}\textbf{x}_{1}^{(0)}&\textbf{x}_{2}^{(0)}&\dots&\textbf{x}_{n}^{(0)}\end{bmatrix}\in\mathbb{C}^{m\times n}$.
while $\max_{j}\lVert\textbf{A}\textbf{x}^{(i+1)}_{j}-\epsilon_{j}^{(i+1)}\textbf{B}\textbf{x}^{(i+1)}_{j}\rVert\,\geq\tau$ do
  Step 1. Chebyshev filtered subspace construction: Compute $\textbf{Y}_{p}^{(i)}=C_{p}(\textbf{H})\textbf{X}^{(i)}$ using the recurrence relation eq. 2 (see algorithm 2).
  Step 2. Rayleigh–Ritz projection: Solve ${\textbf{Y}_{p}^{(i)}}^{\dagger}\textbf{A}\textbf{Y}_{p}^{(i)}\textbf{E}^{(i+1)}={\textbf{Y}_{p}^{(i)}}^{\dagger}\textbf{B}\textbf{Y}_{p}^{(i)}\textbf{E}^{(i+1)}\boldsymbol{\Lambda}^{(i+1)}$ for the Ritz values $\boldsymbol{\Lambda}^{(i+1)}$ and Ritz vectors $\textbf{E}^{(i+1)}$.
  Step 3. Subspace rotation: Set $\textbf{X}^{(i+1)}\leftarrow\textbf{Y}_{p}^{(i)}\textbf{E}^{(i+1)}$.
@@ 2.1.1 Convergence analysis of ChFSI
To ensure completeness, we now provide a mathematical justification for the convergence of ChFSI. The convergence of subspace iteration accelerated by polynomial filters, including Chebyshev polynomials, has been studied in the context of general subspace iteration methods (see, e.g., [46, 38]). The analysis presented here follows the framework of Watkins [46], adapted to explicitly track the maximum principal angle between the Chebyshev polynomial filtered subspace and the target eigenspace, and extended in section 3 to account for approximation errors in the matrix–vector products—a setting not addressed in prior works. Note that $\lVert\cdot\rVert\,$ implies the 2-norm for vectors and vector-induced matrix 2-norm for matrices (spectral norm) throughout this work. In order to analyze convergence, we first define the maximum principal angle between two subspaces [4, 55, 46].
@@
Let $\mathcal{X},\mathcal{Y}\subset\mathbb{C}^{m}$ be subspaces of dimensions $p$ and $q$. Let $\textbf{X}\in\mathbb{C}^{m\times p}$ have orthonormal columns spanning $\mathcal{X}$ and $\textbf{X}_{\perp}\in\mathbb{C}^{m\times(m-p)}$ have orthonormal columns spanning $\mathcal{X}^{\perp}$. Let $\textbf{Y}\in\mathbb{C}^{m\times q}$ have orthogonal columns spanning $\mathcal{Y}$. The maximum principal angle $\angle(\mathcal{X},\mathcal{Y})\in[0,\tfrac{\pi}{2}]$ is defined by
$\tan\angle(\mathcal{X},\mathcal{Y})=\begin{cases}\lVert\textbf{X}_{\perp}^{\dagger}\textbf{Y}\left(\textbf{X}^{\dagger}\textbf{Y}\right)^{+}\rVert\,,&\text{if }\mathrm{rank}(\textbf{X}^{\dagger}\textbf{Y})=q,\\[6.0pt] \frac{\pi}{2},&\text{otherwise},\end{cases}$
where $\dagger$ denotes the conjugate transpose, ++ the Moore–Penrose inverse.
@@
Note that $\hat{\textbf{X}}^{(i)}$ is an orthogonal matrix by virtue of implicit B orthogonalization of $\textbf{X}^{(i)}$ in the Rayleigh-Ritz step. Define the $n\times n$ matrix $\hat{\textbf{Z}}_{1}^{(i)}=\hat{\textbf{U}}_{1}^{\dagger}\hat{\textbf{X}}^{(i)}$ and the $(m-n)\times n$ matrix $\hat{\textbf{Z}}_{2}^{(i)}=\hat{\textbf{U}}_{2}^{\dagger}\hat{\textbf{X}}^{(i)}$. Partitioning the matrix $\hat{\textbf{U}}^{\dagger}\hat{\textbf{X}}^{(i)}$, we have $\hat{\textbf{U}}^{\dagger}\hat{\textbf{X}}^{(i)}=\Bigl[\begin{smallmatrix}\hat{\textbf{Z}}_{1}^{(i)}\\ \hat{\textbf{Z}}_{2}^{(i)}\end{smallmatrix}\Bigr]$. We note that the assumption $\mathcal{S}^{(i)}\cap\mathcal{S}_{\perp}=\left\{0\right\}$ ensures that $\hat{\textbf{Z}}_{1}^{(i)}$ is invertible. This assumption simply requires that the current trial subspace $\mathcal{S}^{(i)}$ is not entirely contained in the unwanted eigenspace; in other words, the trial subspace must retain a nonzero component along every direction of the target eigenspace $\mathcal{S}$. For a randomly initialized subspace this holds with probability one, and the Chebyshev filter strengthens the wanted component at each iteration, so the condition is maintained throughout the algorithm. Using the maximum principal angle given by definition 2.1, we have $\tan{\angle(\mathcal{S}^{(i)},\mathcal{S})}=\lVert\hat{\textbf{Z}}_{2}^{(i)}{{}\hat{\textbf{Z}}_{1}^{(i)}}^{+}\rVert\,$. We can now write
$\mathcal{S}^{(i+1)}=\mathcal{R}(\hat{\textbf{X}}^{(i+1)})=\mathcal{R}(C_{p}(\hat{\textbf{H}})\hat{\textbf{X}}^{(i)}\textbf{E}^{(i+1)})$
We can now write $\hat{\textbf{Z}}_{1}^{(i+1)}=\hat{\textbf{U}}_{1}^{\dagger}\hat{\textbf{X}}^{(i+1)}$ and the $(m-n)\times n$ matrix $\hat{\textbf{Z}}_{2}^{(i+1)}=\hat{\textbf{U}}_{2}^{\dagger}\hat{\textbf{X}}^{(i+1)}$
$\tan{\angle(\mathcal{S}^{(i+1)},\mathcal{S})}\leq$ $\lVert C_{p}(\boldsymbol{\Lambda}_{2})\rVert\,\lVert\hat{\textbf{Z}}_{2}^{(i)}{{}\hat{\textbf{Z}}_{1}^{(i)}}^{+}\rVert\,\lVert C_{p}(\boldsymbol{\Lambda}_{1})^{-1}\rVert\,=\frac{\max_{j\geq n+1}\lvert C_{p}(\lambda_{j})\rvert\,}{\min_{j\leq n}\lvert C_{p}(\lambda_{j})\rvert\,}\tan{\angle(\mathcal{S}^{(i)},\mathcal{S})}$
$=$ $\left\lvert\frac{C_{p}(\lambda_{n+1})}{C_{p}(\lambda_{n})}\right\rvert\,\tan{\angle(\mathcal{S}^{(i)},\mathcal{S})}$
@@ 3 Approximate Chebyshev Filtered Subspace Construction
We note that the computationally dominant step in Chebyshev filtered subspace construction is the evaluation of the matrix multi-vector product $\textbf{H}\textbf{Y}_{k}^{(i)}$ in eq. 2. A straightforward way to accelerate this step is to use approximations in the computation of $\textbf{H}\textbf{Y}_{k}^{(i)}$, allowing for improved efficiency. This should, in principle, allow for the use of various efficient approximate matrix multiplication techniques, including but not limited to mixed-precision arithmetic. However, we note that this requires an understanding of the convergence properties of ChFSI when such approximations are employed. We now adapt theorem 2.2 for the case where approximations are used in computing $\textbf{H}\textbf{Y}_{k}^{(i)}$.